Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 618     Entry time: 04 Jun 2025, 09:44
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 04.06.2025 
General Announcement:
-	 

CTRL:
- Nothing to report.

DA:
- Nothing to report.

EEE:
- Nothing to be reported;
- Yesterday Patrick G. received invitation for the meeting on the 50Hz tests on 17.06 taking place at Desy. Patrick G. will forward it to xo-info.

VAC:
- Not present.

XPD:
- Nothing to report.

DET:
- Nothing to report;
- DSSC issue faced last week: problem was solved; S. Hauf said the follow up was taking place Friday afternoon together with SCS by debugging the detector.

ITDM: 
- Nothing to report.
- A. Galler saw issues while logging in Fast-X and will discuss with J. Szuba on it.

XRO:
- No issues to report.

LAS:
- Not present.
- LIMP activities: instruments that requested use of PP-laser during LIMP shall be secured. Otherwise please contact XO.

UNSYS:
- Positioning HXRSS crystal showed a lost communication in DOOCS server: solved;
- PS in SA2 cell10 was stuck yesterday and is also solved.

SPB:
- Nothing to report.

FXE:
- Not present.

MID:
- Nothing to report.

HED:
- Everything seems to be fine at HED.

SCS:
- Not present.

SQS:
- Nothing to report.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00